This is some Text. This is also some Text which you can’t see in the DVI view. Is it a Lyx or a latex bug ?

Figure 1: Übertragungsfunktion Bandpass Figure 2: Übertragungsfunktion Tiefpass Figure 3: Übertragungsfunktion Hochpass